The specific heats of some elements at 25°C are as follows: aluminum = 0.215 cal/g⋅°C; carbon (graphite) = 0.170 cal/g⋅°C; iron = 0.107 cal/g⋅°C; mercury = 0.0331 cal/g⋅°C.

a) Which element would require the smallest amount of heat to raise the temperature of 100. g of the element by 10.°C?

b) If the same amount of heat needed to raise the temperature of 1 g of aluminum by 59°C were applied to 1 g of mercury, by how many degrees would its temperature be raised? deltaT Hg= ............... Celises

c) If a certain amount of heat is used to raise the temperature of 1.6 g of iron by 10.°C, the temperature of 1 g of which element would also be raised by 10.°C, using the same amount of heat?

-aluminum

-carbon (graphite)

-mercury



1
Expert's answer
2021-01-20T03:16:21-0500

1) Element with the smallest specific heat would require the smallest amount of heat. Answer: mercury.

2) 0.215*59 = 0.0331*x, deltaT = x = 383.2°C

3) 1.6*0.107*10 = 1*x*10, x = 0.1712. Answer: carbon.